• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name rousseau

    Log entry time 08:01:42 on August 03, 2006

    Entry number 177074

    keyword=shift summary

    target LH2, radiator out, beam 20 microA run 2042 in progress beam 14 microA, radiator moved in 4% run 2043-2052 : LH2 problem during the run 2048 : the alarm is red for the lecroy hv (beamline), problem fixed by turnin the hv off and on again

    spectrometer moved to 18.84, momentum set to 0.5421 (new values for current, see halog #177071) target moved to optics radiator out spectrometer motion done... beam up... we start the run 2053 (optics) and left...

    bye bye see you